OpenStudy (anonymous):

Find cos2(theta) if cos(theta) =3/4. Do I need to create a triangle, solve for sin, then set it equal to 1-2sin^2(theta)?

OpenStudy (anonymous):

Or do I need to us a double angle identity?

OpenStudy (dumbcow):

either way will work \[\cos 2\theta = 2 \cos^2 \theta - 1\]

OpenStudy (dumbcow):

i guess its simpler to just use double angle identity here

OpenStudy (anonymous):

hmm...the double and is \[\cos^2\theta-\sin^2\theta\] right?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

sorry angle

OpenStudy (anonymous):

so since the \[\cos = \frac{ 3 }{ 4 }\] we can replace cos()theta^2 with (3/4)^2?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

but what happens to the sin^2(theta)?

OpenStudy (dumbcow):

\[\sin^2 + \cos^2 = 1\] make the substitution \[\sin^2 = 1 - \cos^2\]

OpenStudy (anonymous):

that's what I thought

OpenStudy (anonymous):

So its going to be\[\frac{ 9 }{ 16 }-(1-\cos^2(\theta) ?\]

OpenStudy (anonymous):

Got it! Answer is 1/8.